Qd1 = 90 – 2P1 + 3P2 – 5P3 ; Qs1 = P1 – 10

Qd2 = 36 + 3P1 – 3P2 + 2P3 ; Qs2 = –14 + P2

Qd3 = 45 – 3P1 + 3P2 – 3P3 ; Qs3 = P3 – 20

a. Determine the equilibrium prices and quantities for the three commodity Market

model.


1
Expert's answer
2022-01-17T12:37:17-0500

At Equilibrium, Quantity demanded = Quantity supplied

"Q_{d1}=Q_{S2}"

"90-2P_1+ 3P_2-5P_3= P_1-10"

"=100-3P_1+ 3P_2-5P_3= 0" ....................(i)

"Q_{d2}=Q_{s2}"


"36+ 3P_1-3P_2+2P_3=-14+P_2"

"=50+3P_1-4P_2+2P_3= 0" .........................(ii)


"Q_{d3}=Q_{s3}"

"45-3P_1+ 3P_2-3P_3= P_3-20"

"=25-3P_1+ 3P_2-2P_3= 0" ...........................(iii)

Solve Eqns (I) and (ii) simultaneously by elimination method by adding the two eqns

"100-3P_1+ 3P_2-5P_3= 0"

"50+3P_1-4P_2+2P_3= 0"

="150-P_2-3P_3=0" ..............................................(iv)

Solve Eqns (ii) and (iii) simultaneously by adding the two Eqns

"50+3P_1-4P_2+2P_3= 0"

"25-3P_1+ 3P_2-2P_3= 0"

="75-P_2"

"P_2= 75"

From Eqn (iv),

"150-75-3P_3=0"

"3P_3=75"

"P_3= 25"

From Eqn (I)

"100-3P_1+ 3(75)-5(25)= 0"

"3P_1= 200"

"P_1= \frac{200}{3}= 66.7"

"Q_1= 66.7-10= 56.7"

"Q_2= -14+75= 61"

"Q_3= 25-20=5"